# An Interdisciplinary and Cross-Task Review on Missing Data Imputation

### v1

@@ 8.2 Exact Recovery Guarantee
Candès and Recht (2009) proved that a rank-$r$ matrix of size $n\times n$ can be recovered exactly via nuclear norm minimization when the number of observed entries (sampled uniformly at random) is larger than $Cn^{1.2}r\log n$ for some positive numerical constant $C$. The bound was further extended or improved by a few studies such as (Candes and Plan, 2010; Chen et al., 2014). Regarding the low-rank factorization model for matrix completion, nonconvex but more efficient than nuclear norm minimization, a few scholars proposed alternating minimization and gradient descent algorithms that can exactly recover low-rank matrices (Keshavan, 2012; Jain et al., 2013; Hardt, 2014; Sun and Luo, 2015). Compared to low-rank matrix completion, exact recovery guarantees for low-rank tensor completion is more difficult to obtain. Nevertheless, there have been a few guaranteed algorithms for tensor completion under CP decomposition (Krishnamurthy and Singh, 2013; Mu et al., 2014; Jain and Oh, 2014; Potechin and Steurer, 2017; Liu and Moitra, 2020), Tucker decomposition (Huang et al., 2014), and t-SVD (Zhang and Aeron, 2016).
@@ 8.3 Generalization Error Bound
Srebro and Shraibman (2005) analyzed the generalization error bounds of rank, nuclear norm, and max-norm constrained matrix completion methods under a uniform sampling distribution. Shamir and Shalev-Shwartz (2014) provided a generalization bound for matrix completion with nuclear norm minimization under a milder assumption. (Negahban and Wainwright, 2012; Fan et al., 2019) analyzed the generalization bound for Schatten-$p$ quasi-norm based matrix completion. Fan (2022a) provided generalization bounds for deep matrix and tensor factorization models on incomplete data. Fan et al. (2023) analyzed the generalization bounds of Schatten-$p$ norm minimization for LRTC and showed that for LRTC with Schatten-$p$ quasi-norm regularizer on $d$-order tensors, $p=1/d$ is always better than any $p>1/d$ in terms of the generalization ability. Wu and Fan (2024) analyzed the generalization bound of total-variation regularized tensor completion. Ledent and Alves (2024) provided a tighter bound for matrix completion with Schatten-$p$ quasi-norm constraint.
@@ 9.1 Benchmarks and Toolboxes
Despite the extensive literature on missing data imputation, comprehensive toolboxes and benchmarks remain relatively limited. Several studies have sought to fill this gap. For instance, Folch-Fortuny et al. (2016) provided a MATLAB toolbox for missing data imputation and the imputation methods are mainly based on PCA and linear regression. Josse and Husson (2016) developed an R package for handling missing data. It includes principal component analysis for continuous variables, multiple correspondence analysis for categorical variables, factorial analysis on mixed data for both continuous and categorical variables, and multiple factor analysis for multi-table data. Rubinsteyn and Feldman (2020) developed fancyimpute, a Python toolbox of missing data imputation, that included simple imputation, kNN, and a few LRMC methods. Woźnica and Biecek (2020) evaluated the empirical effectiveness of 7 imputation algorithms for 5 predictive models (logistic regression, classification tree, random forest, etc.) on 13 datasets. Jäger et al. (2021) compared 6 imputation methods on 69 datasets under MCAR, MAR, and MNAR missingness patterns (with different missing rates) and evaluated by regression and classification downstream tasks. Miao et al. (2022) experimentally compared 19 imputation algorithms on 15 real-world benchmark datasets. Perez-Lebel et al. (2022) compared 6 imputation methods in 10 classification and 3 regression tasks on 4 publicly available health databases. Pereira et al. (2024) tested 6 imputation methods in an experimental setup that covers 10 datasets and 5 missingness levels (10% to 80%) in different MNAR settings. Empirically, they found that, for most missing rates and datasets, multiple imputation by chained equations performed best, whereas autoencoders showed promising results at higher missingness rates. Cabrera-Sánchez et al. (2024) provided a comparison of classical and deep learning imputation methods on 21 heterogeneous datasets in various areas. Pons-Suñer et al. (2024) developed a toolbox to assess the reliability of various imputation methods, select the best imputer for any feature or group of features, and filter out features that do not meet quality criteria. Toye et al. (2025) evaluated 12 imputation methods on two time-series datasets only. Richter et al. (2025) proposed ImputeBenc, which includes 4 imputation methods only, though more methods can be inserted into the workflow. There are more benchmarking papers on missing data imputation, most of which focus on specific data types or problems. For convenience, we summarize those published after 2020 in Table 6.
@@
(Woźnica and Biecek, 2020) general 7 13
Jäger et al. (2021) general 6 69
Ahn et al. (2022) time series 6 4
(Miao et al., 2022) general 19 15
(Perez-Lebel et al., 2022) healthcare 6 4
(Sun et al., 2023) general 6 10
Kazijevs and Samad (2023) time series 7 5
(Pereira et al., 2024) general 6 10
(Cabrera-Sánchez et al., 2024) general 8 21
(Gama et al., 2024) geological 8 3
(Prakash et al., 2024) health survey 4 10
(Gendre et al., 2024) biological 6 2
(Du et al., 2024b) time series 28 8
(Toye et al., 2025) time series 12 2
(Zhang et al., 2025c) healthcare 5 3
(Richter et al., 2025) general 4
@@ 9.2 Evaluation Metrics
The second group is based on distance metrics between distributions such as the Wasserstein distance (WD) (Villani et al., 2008) and maximum mean discrepancy (MMD) (Gretton et al., 2012). Let $\mathbf{X}\in\mathbb{R}^{d\times n_{x}}$ and $\mathbf{Y}\in\mathbb{R}^{d\times n_{y}}$ be two discrete distributions in $\mathbb{R}^{d}$, WD2 (Wasserstein-2 distance) is calculated by
@@
As can be seen, the computations of WD and MMD are much more time-consuming than the imputation errors, but they can be applied to cases where the ground truth of the missing values is not available. For example, let $\tilde{\mathbf{X}}_{a}$ be an incomplete data matrix, for which the ground truth of the missing values is not available. Let $\hat{\mathbf{X}}_{a}$ be the imputed data matrix given by some imputation method. If there is another complete data matrix ${\mathbf{X}}_{b}$ from the same distribution as $\mathbf{X}_{a}$, we can measure the WD between $\hat{\mathbf{X}}_{a}$ and $\mathbf{X}_{b}$, where a smaller WD means a better imputation performance. WD has been utilized in a few papers such as (Muzellec et al., 2020; Jarrett et al., 2022).
The third group of evaluation metrics assesses performance on downstream tasks—specifically regression, classification, and clustering. For regression, the most common measures are the Mean Absolute Error MAE and RMSE. For classification, standard metrics include accuracy, precision, recall, F1-score, AUROC, and AUPRC; Grandini et al. (2020) provided a comprehensive survey of these measures. Clustering metrics fall into two families: internal indices, which do not require ground-truth labels (e.g., the Davies–Bouldin index and Silhouette coefficient), and external indices, which do (e.g., clustering accuracy, normalized mutual information (NMI), and the adjusted rand index (ARI)).
@@ 10.1 Model Selection and Hyperparameter Tuning for Missing Data Imputation
Method or model selection for missing data imputation is challenging. Almost every imputation method has at least one hyperparameter and the best values for the hyperparameters vary across different datasets. For each dataset, we need to split it into two subsets, one for training and the other for validation, to find the best hyperparameters. This will be time-consuming when the number of hyperparameters is large and the imputation method has a high-time complexity. For instance, for a deep learning-based imputation method, we need to determine the number of layers of the neural network, the width of each layer, the learning rate, the number of iterations, the coefficient of weight decay, and algorithm-specific hyperparameters; if each has 4 candidate values, we need to use grid search over at least 4096 combinations, meaning training 4096 neural networks, which is extremely costly when the dataset is large. To accelerate the hyperparameter tuning, we may use more effective strategies such as Bayesian optimization (Snoek et al., 2012) or other automated machine learning (AutoML) techniques (Hutter et al., 2019) instead of grid search. Although some works, such as (Jarrett et al., 2022), have used AutoML to tune the hyperparameters of imputation methods, the related studies are limited, and there is a large room for improvement. Future work may consider the following points:
@@ 10.2 Privacy Protection in Missing Data Imputation
Data privacy is the cornerstone of trustworthy data science: without it, individuals can be re-identified from supposedly anonymized models, exposing medical records, financial details, or location histories. For missing data imputation, we also need to consider the problems of data leakage and privacy protection. Privacy-preserving techniques like differential privacy (Dwork, 2006) and federated learning (Kairouz et al., 2021) enable organizations to unlock collective insights across silos—think hospitals training shared cancer-detection models without ever moving patient data—turning privacy from a legal checkbox into a competitive advantage. In federated-learning scenarios, the training data are distributed across multiple clients, and any sharing or transmission between clients—or between a client and the central server, if one exists—is forbidden, thereby protecting data privacy to some extent. This protection is often further strengthened by differential privacy, with a trade-off between accuracy and privacy.
Recently, there have been a few studies considering federated learning for missing data imputation (Li et al., 2021; Abbasi et al., 2023; Lian et al., 2025; Fang et al., 2025a; Li et al., 2025c), but there are still a few challenges.
  • It is difficult to extend nonparametric imputation methods such as kNN and tree-based methods to federated learning scenarios (Zhang et al., 2024a; Li et al., 2020). For example, regarding kNN, we cannot directly compute the distance between two data points located in different clients.
@@ 10.4 Pretrained and Universal Models for Missing Data Imputation
In recent years, pretrained models, such as BERT (Devlin et al., 2019), CLIP (Radford et al., 2021), and GPTs (Achiam et al., 2023), have become the centerpiece of AI research, delivering breakthrough performance on a wide spectrum of real-world problems. A pretrained model is a large neural network whose parameters have already been learned from massive, general-domain data—text, images, audio, protein sequences, or code—so it arrives packed with broad patterns and world knowledge; instead of training from scratch, practitioners simply fine-tune or prompt it with a handful of examples, slashing data, compute, and time by orders of magnitude. It is natural to consider using pretrained models in missing data imputation or even establishing a universal model for missing data imputation. Here, a universal model for missing data imputation refers to one that, once well trained, can impute missing values for any dataset—or at least a broad class of datasets from different domains—without requiring further tuning or with only light fine-tuning.
As discussed in Section 5.7, a few scholars have explored the use of LLMs to assist in missing data imputation (Huang and Hao, 2024; Hayat and Hasan, 2024; Fang et al., 2025b) and achieved some inspiring results in general missing data imputation tasks. However, there are still limitations:
@@
  • Since existing methods heavily the textualization strategy for the tabular data or other data types, it is important to invent more elegant textualization strategies. It is also interesting to use LLM to improve missing data imputation without textualization.

### v2

@@ 8.2 Exact Recovery Guarantee
CandesRecht2009 proved that a rank-$r$ matrix of size $n\times n$ can be recovered exactly via nuclear norm minimization when the number of observed entries (sampled uniformly at random) is larger than $Cn^{1.2}r\log n$ for some positive numerical constant $C$. The bound was further extended or improved by a few studies such as (candes2010matrix; MC_icml2014c1_chenc14). Regarding the low-rank factorization model for matrix completion, nonconvex but more efficient than nuclear norm minimization, a few scholars proposed alternating minimization and gradient descent algorithms that can exactly recover low-rank matrices (keshavan2012efficient; jain2013low; hardt2014understanding; ruoyu15). Compared to low-rank matrix completion, exact recovery guarantees for low-rank tensor completion is more difficult to obtain. Nevertheless, there have been a few guaranteed algorithms for tensor completion under CP decomposition (krishnamurthy2013low; mu2014square; jain2014provable; pmlr-v65-potechin17a; NEURIPS2020_dab1263d), Tucker decomposition (huang2014provable), and t-SVD (zhang2016exact).
@@ 8.3 Generalization Error Bound
srebro2005rank analyzed the generalization error bounds of rank, nuclear norm, and max-norm constrained matrix completion methods under a uniform sampling distribution. shamir14a provided a generalization bound for matrix completion with nuclear norm minimization under a milder assumption. (negahban2012restricted; fan2019factor) analyzed the recovery error bound for Schatten-$p$ quasi-norm based matrix completion. fan2021multi provided generalization bounds for deep matrix and tensor factorization models on incomplete data. fan2023euclideannorminduced analyzed the generalization bounds of Schatten-$p$ norm minimization for LRTC and showed that for LRTC with Schatten-$p$ quasi-norm regularizer on $d$-order tensors, $p=1/d$ is always better than any $p>1/d$ in terms of the generalization ability. wu2024smooth analyzed the generalization bound of total-variation regularized tensor completion. ledent2024generalization provided a tighter bound for matrix completion with Schatten-$p$ quasi-norm constraint.
@@ 9.1 Benchmarks and Toolboxes
Despite the extensive literature on missing data imputation, comprehensive toolboxes and benchmarks remain relatively limited. Several studies have sought to fill this gap. For instance, folch2016missing provided a MATLAB toolbox for missing data imputation and the imputation methods are mainly based on PCA and linear regression. josse2016missmda developed an R package for handling missing data. It includes principal component analysis for continuous variables, multiple correspondence analysis for categorical variables, factorial analysis on mixed data for both continuous and categorical variables, and multiple factor analysis for multi-table data. fancyimpute developed fancyimpute, a Python toolbox of missing data imputation, that included simple imputation, kNN, and a few LRMC methods. woznica2020does evaluated the empirical effectiveness of 7 imputation algorithms for 5 predictive models (logistic regression, classification tree, random forest, etc.) on 13 datasets. jager2021benchmark compared 6 imputation methods on 69 datasets under MCAR, MAR, and MNAR missingness patterns (with different missing rates) and evaluated by regression and classification downstream tasks. miao2022experimental experimentally compared 19 imputation algorithms on 15 real-world benchmark datasets. perez2022benchmarking compared 6 imputation methods in 10 classification and 3 regression tasks on 4 publicly available health databases. pereira2024imputation tested 6 imputation methods in an experimental setup that covers 10 datasets and 5 missingness levels (10% to 80%) in different MNAR settings. Empirically, they found that, for most missing rates and datasets, multiple imputation by chained equations performed best, whereas autoencoders showed promising results at higher missingness rates. cabrera2024benchmark provided a comparison of classical and deep learning imputation methods on 21 heterogeneous datasets in various areas. pons2024iti developed a toolbox to assess the reliability of various imputation methods, select the best imputer for any feature or group of features, and filter out features that do not meet quality criteria. toye2025benchmarking evaluated 12 imputation methods on two time-series datasets only. richter2025imputebench proposed ImputeBenc, which includes 4 imputation methods only, though more methods can be inserted into the workflow. There are more benchmarking papers on missing data imputation, most of which focus on specific data types or problems. For convenience, we summarize those published after 2020 in Table 6.
@@
(woznica2020does) general 7 13
(jager2021benchmark) general 6 69
(ahn2022comparison) time series 6 4
(miao2022experimental) general 19 15
(perez2022benchmarking) healthcare 6 4
(sun2023deep) general 6 10
kazijevs2023deep time series 7 5
(pereira2024imputation) general 6 10
(cabrera2024benchmark) general 8 21
(gama2024imputation) geological 8 3
(prakash2024benchmarking) health survey 4 10
(gendre2024benchmarking) biological 6 2
(du2024tsi) time series 28 8
(toye2025benchmarking) time series 12 2
(zhang2025comparison) healthcare 5 3
(richter2025imputebench) general 4
@@ 9.2 Evaluation Metrics
The second group is based on distance metrics between distributions such as the Wasserstein distance (WD) (villani2008optimal) and maximum mean discrepancy (MMD) (gretton2012kernel). Let $\mathbf{X}\in\mathbb{R}^{d\times n_{x}}$ and $\mathbf{Y}\in\mathbb{R}^{d\times n_{y}}$ be two discrete distributions in $\mathbb{R}^{d}$, WD2 (Wasserstein-2 distance) is calculated by
@@
As can be seen, the computations of WD and MMD are much more time-consuming than the imputation errors, but they can be applied to cases where the ground truth of the missing values is not available. For example, let $\tilde{\mathbf{X}}_{a}$ be an incomplete data matrix, for which the ground truth of the missing values is not available. Let $\hat{\mathbf{X}}_{a}$ be the imputed data matrix given by some imputation method. If there is another complete data matrix ${\mathbf{X}}_{b}$ from the same distribution as $\mathbf{X}_{a}$, we can measure the WD between $\hat{\mathbf{X}}_{a}$ and $\mathbf{X}_{b}$, where a smaller WD means a better imputation performance. WD has been utilized in a few papers such as (muzellec2020missing; jarrett2022hyperimpute).
The third group of evaluation metrics assesses performance on downstream tasks—specifically regression, classification, and clustering. For regression, the most common measures are the Mean Absolute Error MAE and RMSE. For classification, standard metrics include accuracy, precision, recall, F1-score, AUROC, and AUPRC; grandini2020metrics provided a comprehensive survey of these measures. Clustering metrics fall into two families: internal indices, which do not require ground-truth labels (e.g., the Davies–Bouldin index and Silhouette coefficient), and external indices, which do (e.g., clustering accuracy, normalized mutual information (NMI), and the adjusted rand index (ARI)).
@@ 10.1 Model Selection and Hyperparameter Tuning for Missing Data Imputation
Method or model selection for missing data imputation is challenging. Almost every imputation method has at least one hyperparameter and the best values for the hyperparameters vary across different datasets. For each dataset, we need to split it into two subsets, one for training and the other for validation, to find the best hyperparameters. This will be time-consuming when the number of hyperparameters is large and the imputation method has a high-time complexity. For instance, for a deep learning-based imputation method, we need to determine the number of layers of the neural network, the width of each layer, the learning rate, the number of iterations, the coefficient of weight decay, and algorithm-specific hyperparameters; if each has 4 candidate values, we need to use grid search over at least 4096 combinations, meaning training 4096 neural networks, which is extremely costly when the dataset is large. To accelerate the hyperparameter tuning, we may use more effective strategies such as Bayesian optimization (snoek2012practical) or other automated machine learning (AutoML) techniques (hutter2019automated) instead of grid search. Although some works, such as (jarrett2022hyperimpute), have used AutoML to tune the hyperparameters of imputation methods, the related studies are limited, and there is a large room for improvement. Future work may consider the following points:
@@ 10.2 Privacy Protection in Missing Data Imputation
Data privacy is the cornerstone of trustworthy data science: without it, individuals can be re-identified from supposedly anonymized models, exposing medical records, financial details, or location histories. For missing data imputation, we also need to consider the problems of data leakage and privacy protection. Privacy-preserving techniques like differential privacy (dwork2006differential) and federated learning (kairouz2021advances) enable organizations to unlock collective insights across silos—think hospitals training shared cancer-detection models without ever moving patient data—turning privacy from a legal checkbox into a competitive advantage. In federated-learning scenarios, the training data are distributed across multiple clients, and any sharing or transmission between clients—or between a client and the central server, if one exists—is forbidden, thereby protecting data privacy to some extent. This protection is often further strengthened by differential privacy, with a trade-off between accuracy and privacy.
Recently, there have been a few studies considering federated learning for missing data imputation (10.14778/3503585.3503598; abbasi2023fast; lian2025federated; fang2025msfgan; 10.1016/j.eswa.2025.126543), but there are still a few challenges.
  • It is difficult to extend nonparametric imputation methods such as kNN and tree-based methods to federated learning scenarios (zhang2024fedknn; Li_Wen_He_2020). For example, regarding kNN, we cannot directly compute the distance between two data points located on different clients.
@@ 10.4 Pretrained and Universal Models for Missing Data Imputation
In recent years, pretrained models, such as BERT (devlin2019bert), CLIP (radford2021learning), and GPTs (achiam2023gpt), have become the centerpiece of AI research, delivering breakthrough performance on a wide spectrum of real-world problems. A pretrained model is a large neural network whose parameters have already been learned from massive, general-domain data—text, images, audio, protein sequences, or code—so it arrives packed with broad patterns and world knowledge; instead of training from scratch, practitioners simply fine-tune or prompt it with a handful of examples, slashing data, compute, and time by orders of magnitude. It is natural to consider using pretrained models in missing data imputation or even establishing a universal model for missing data imputation. Here, a universal model for missing data imputation refers to one that, once well trained, can impute missing values for any dataset—or at least a broad class of datasets from different domains—without requiring further tuning or with only light fine-tuning.
As discussed in Section 5.7, a few scholars have explored the use of LLMs to assist in missing data imputation (huang2024missing; hayat2024context; fang2025spatiotemporal) and achieved some inspiring results in general missing data imputation tasks. However, there are still limitations:
@@
  • Since existing methods heavily rely on the textualization strategy for the tabular data or other data types, it is important to invent more elegant textualization strategies. It is also interesting to use LLM to improve missing data imputation without textualization.